The following variable production costs apply to goods made by Baird Manufacturing Corporation: Item Materials Labor Variable1
Answer #1
13000 23000 33000
Total variable cost (11.75*13000)=$152750 (11.75*23000)=$270250 (11.75*33000)=$387750

NOTE:Total fixed costs and variable cost per unit do not change with change in units